ROTOR-ROTOR SPECTRA AND DYNAMICS: COUPLING BETWEEN EXTREMES IN INTERACTION STRENGTH AND SPECTRAL COMPLEXITY

\underline{WILLIAM G. HARTER} and HORACE T. CROGMAN, Department of Physics University of Arkansas Fayetteville, AR 72701.

Buckminsterfullerene (C_{60}) is the roundest commercially available deformable spherical rotors (DSR),^{a}. The ”Buckyballs” are unique in that they have tri-axial rotation even in the fullerite solid-state above 140K; a box of frictionless ball bearings! Also, the C_{60} cage is large enough to hold atoms and maybe even small molecules. In gas-phase this opens the possibility of a fully triaxial concentric DSR-DST rotor-rotor as well as many of the other more theoretically tractable cases like DSR-RDR or RST-RUP listed in the preceding talk. (C_{60} solid-phase is a hetrocentric (DSR)^{N} fcc crystal.) Finally, C_{60} is practically unique in that it comes in distinct isotopic varieties, most notably a ”Bose-ball” ^{12}C_{60}, a ”mixed-ball” ^{13}C_{59} and a ”Fermi-ball” of (^{13}C_{60}).^{b}, ^{c} Differences in symmetry between these casesis enormous. Bose-exclusion kills all but the Ag rovibrational species of ^{12}C_{60} and reduces spectral congestion by a factor of 60. In contrast, Pauli-Fermi-exclusion assigns 2^{60} 1.5E18 hyperfine levels to each rovibrational level of ^{13}C_{60}. Mixed symmetry fullerenes ^{13}C_{m}-^{12}C_{60−m}[3] lie between the two extremes. The first ”mixed-ball” ^{13}C_{60} - ^{12}C_{59} corresponds to a strong-coupling limit of an hcDSR-RDR model. This observation helps to clarify the enormous symmetry-exclusion diversity. At the bottom of the rotor-rotor symmetry hierarchy are the hetrocentric hcRDR-RDR or hcRAT-RAT arrangements of elongated molecules connected more or less loosely end-to-end, that is, links of folding polymer chains. Such chains and links are nano-sized quantum mechanical analogs of ancient slings, whips, and trebuchets with a devilish dynamics to match! And, biological systems cannot live without them.

aH. W. Kroto, J. R. Heath, S. C. O’Brien, R. F. Curl, and R. E. Smalley,Nature 318, 162(1985), W. Kratschmer, et. al.

Molecules (cf. SiF_{4})

bW. G. Harter and D. E. Weeks, Chem. Phys. Letters 132, 387 (1986), J. Chem. Phys. 90, 4744 (1989).

cW. G. Harter and T. C. Reimer,Chem. Phys. Letters 194, 230 (1992), J. Chem. Phys. 106, 1326 (1997).